# Positive microlocal holonomies are globally regular

### v1

@@ 2.3. A $\mathfrak{t}$-action on the derived moduli stacks of sheaves
For the case of $\mathfrak{t}$ with multiple basepoints in one component of $\Lambda$, we reduce to the case of one basepoint per component as follows. Consider a subset $\mathfrak{t}^{\prime}\subset\mathfrak{t}\subset\Lambda$ that still has at least one point per component. The homotopy colimit diagram (2.2) can be broken down into the following diagram
[…]
@@
${\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})}$${\mathfrak{M}_{n}({\Lambda,\mathfrak{t}^{\prime}})}$${\mathfrak{M}_{n}({\Lambda})}$${\operatorname{BGL}_{n}}$${\displaystyle\prod_{i=1}^{|\mathfrak{t}|-|\mathfrak{t}^{\prime}|}\operatorname{BGL}_{n}\times\operatorname{BGL}_{n}}$${\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{BGL}_{n}.}$
@@
Remark 2.7.
The case of microlocal rank 1, in which this action factors through the Abelianization of $F_{p}$, is of particular interest to study commutative cluster algebras, and we will momentarily focus on it. Indeed, by Example 2.5 (and Lemma 2.6), if we restrict to microlocal rank 1 we obtain an action factoring through the monodromies of (microlocal) Abelian local systems. In terms of algebraic geometry, this translates into some of the $\operatorname{GL}_{1}$-actions studied in the literature, see e.g. [CGGS20, Section 2.2]. In the literature, these are often referred to as torus actions because $k=\mathbb{C}$ is chosen and $\operatorname{GL}_{1}(\mathbb{C})\cong\mathbb{C}^{*}$ is an algebraic torus. $\Box$
@@ 2.4. Microlocal rank 1 case of the basepoint action
Let us focus on $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\subset\mathfrak{M}({\Lambda,\mathfrak{t}})$, the component of $\mathfrak{M}({\Lambda,\mathfrak{t}})$ consisting of sheaves singularly supported in $(\Lambda,\mathfrak{t})$ with microlocal rank $1$. This is the component that has been most studied in the literature and it is of particular interest due to its connection to cluster algebras, see e.g. [CG22, CW24, CZ22, CGGS20, CGGS21, CGG+22]. The component $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ parametrizes sheaves with singular support on $\Lambda$ such that their microstalks at the basepoints of $\mathfrak{t}$ are 1-dimensional $k$-modules, all commonly identified. Similarly, we denote by $\mathfrak{M}_{1}({\Lambda})$ the component of $\mathfrak{M}({\Lambda})$ consisting of microlocal rank $1$ sheaves singularly supported in $\Lambda$.
@@
$\mathfrak{M}(\operatorname{\mbox{Mod}}(k))\quad\mbox{and}\quad\displaystyle\mathfrak{M}\left(\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)\right).$
@@
(2.5) ${\mathfrak{M}_{1}(\Lambda,\mathfrak{t})}$${\mathfrak{M}_{1}(\Lambda)}$${BH}$${BG,}$$\scriptstyle{\mathfrak{M}(m_{\Lambda,\mathfrak{t}}^{\ell})}$$\scriptstyle{B\Delta}$
@@
Remark 2.8.
In the general non-Abelian $\operatorname{GL}_{n}$ case, the diagonal action of $H$ is non-trivial and $H$ is not a normal subgroup. In appropriate coordinates, this diagonal subgroup acts via conjugation. In the $n=1$ case this simplifies, as the action becomes trivial, and we can reduce to a $(G/H)$-principal bundle and merely have a factor of $BH$ in the moduli stack. $\Box$
@@ 2.5. A few examples
Here are descriptions of the derived stacks $\mathfrak{M}_{1}({\Lambda})$ and $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ and the basepoint action in some cases of interest.
@@ 2.5.1. Legendrian knots
Let $\Lambda\subset(\mathbb{R}^{3},\xi_{\text{st}})\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{\text{st}})$ be a Legendrian knot with $\mathfrak{t}=\{t\}$ consisting of a single basepoint. Then $\Delta:H\longrightarrow G$ is the identity morphism and the homotopy pullback (2.5) is base change along the identity morphism. Therefore $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\cong\mathfrak{M}_{1}({\Lambda})$ are isomorphic. Two remarks:
@@
    Let $\Lambda\subset(\mathbb{R}^{3},\xi_{\text{st}})$ be a Legendrian knots of the form $\Lambda=\Lambda_{\beta}$, $\beta$ a positive braid and $\Lambda_{\beta}$ its $(-1)$-closure, see [CN22, Section 2.2]. Then the derived stacks $\mathfrak{M}_{1}({\Lambda_{\beta}})$ are isomorphic to quotient stacks $[X_{\beta}/\operatorname{GL}_{1}]$ where $X_{\beta}$ are smooth affine varieties and $\operatorname{GL}_{1}$ acts trivially on $X_{\beta}$, cf. [CW24, Section 4.1]. As stated in (ii) below, these $X_{\beta}$ are closely related to the braid varieties $X(\beta)$ studied in [CGGS20, CGGS21, CGG+22]. In short, in this case we have
@@
    where $X_{\beta}$ is a smooth affine scheme. Also, equations expressing $X_{\beta}$ as a complete intersection can be readily read from a braid word for $\beta$, cf. [CGG+22, Section 3] or [CN22, Section 5].
@@
    For a Legendrian knot with basepoints $\mathfrak{t}=\{t_{1},\ldots,t_{|\mathfrak{t}|}\}$, the action of $G/H$ on $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ is free and $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\cong\mathfrak{M}_{1}({\Lambda})\times(\operatorname{GL}_{1})^{|\mathfrak{t}|-1}$, as per Equation (2.4). Let $\Lambda\subset(\mathbb{R}^{3},\xi_{\text{st}})$ be a Legendrian knot of the form $\Lambda=\Lambda_{\beta\delta(\beta)}$ , with one basepoint per strand of $\beta$. By [CL23, Proposition 5.8], we have
@@
    where $X(\beta)$ is the braid variety associated to $\beta$, see e.g. [CGG+22, Section 3]. Specifically, per Equation (2.4), we have a regular isomorphism $X(\beta)\cong X_{\beta\delta(\beta)}\times\operatorname{GL}_{1}^{n-1}$, where $\delta(\beta)$ is a braid lift of the Demazure product of $\beta$.
@@ 2.5.2. Hopf link, Part I: $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ and $\mathfrak{M}_{1}({\Lambda})$
Let $\Lambda=\Lambda_{1}\cup\Lambda_{2}\subset(\mathbb{R}^{3},\xi_{\text{st}})\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{\text{st}})$ be the max-tb Hopf link, with both components $\Lambda_{1},\Lambda_{2}$ max-tb unknots. Consider the set of basepoints $\mathfrak{t}=\{t_{1},t_{2}\}$, with $t_{i}\in\Lambda_{i}$, one per component. The homotopy pullback (2.5) and the basepoint action are explicitly described as follows, cf. [CW24, Section 4.1]. Consider the polynomial ring $\mathbb{Z}[x,y]$ and its localization $\mathbb{Z}[x,y]_{(1+xy)}$ at the principal ideal $(1+xy)$. Define the affine scheme $X:=\operatorname{Spec}\mathbb{Z}[x,y]_{(1+xy)}$. The basepoint action from Definition 2.4 can be described as follows: consider the algebraic action of $G:=\operatorname{GL}_{1}\times\operatorname{GL}_{1}$ on $X$ defined by
@@
Here we interpret the scheme $X$ as a 0-geometric Artin 0-stack, cf. [TV08, Prop. 2.1.2.1]. Note also that the diagonal subgroup $\Delta(H)=\{(t_{1},t_{2}):t_{1}=t_{2}\}\subset G$, $\Delta:H\longrightarrow G$ the diagonal morphism, acts trivially on $X$. Then the moduli stacks of sheaves appearing in (2.5) are
@@
The two input morphisms $\mathfrak{M}_{1}(m^{\ell}_{\Lambda,\mathfrak{t}}):[X/G]\longrightarrow BG$ and $B\Delta:BH\longrightarrow BG$ in the homotopy pullback (2.5) are given by the canonical map $[X/G]\longrightarrow BG$, induced by $X\longrightarrow\{\mbox{pt}\}$, and the diagonal map $B\Delta:\operatorname{BGL}_{1}\longrightarrow\operatorname{BGL}_{1}\times\operatorname{BGL}_{1}$. The induced morphism $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\longrightarrow BH$ is also given by the canonical map $[X/H]\longrightarrow BH$. The last induced $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\longrightarrow\mathfrak{M}_{1}({\Lambda})$ in (2.5) is an algebraic quotient by $G/H$. This latter quotient can be realized by the action of $K:=\operatorname{GL}_{1}$ on $[X/H]$ induced by the action
@@
where we have taken the slice $\tau=t_{1}t_{2}^{-1}$ to represent the quotient $K\cong G/H$. In summary, with the notation and actions above, the homotopy pullback (2.5) reads
(2.7) ${{[X/H]}}$${{[X/G]}}$${BH}$${BG,}$$\scriptstyle{B\Delta}$
where both vertical maps are the canonical projections induced by $X\longrightarrow\mbox{pt}$ and both horizontal maps can be modeled after $(G/H)$-principal bundles. In this context, the $(G/H)$-principal bundle $[X/H]\longrightarrow[X/G]$ is classified by the composition $[X/G]\longrightarrow BG$ with the natural map $BG\longrightarrow B(G/H)$.
@@ 2.5.3. Hopf link, Part II: Lagrangian fillings
The quotient stack $\mathfrak{M}_{1}({\Lambda})\cong[X/G]$ in Example 2.5.2 has a natural interpretation in terms of Lagrangian fillings of the Hopf link, as follows. There are three types of orbits in $X=\operatorname{Spec}\mathbb{Z}[x,y]_{(1+xy)}$ for the action (2.6):
@@
    If $xy=\alpha$, $\alpha\in k$ and $\alpha\neq 0,-1$, then $K$ acts freely and transitively on $O_{\alpha}:=\{xy=\alpha\}\subset X$. Therefore, the orbit $O_{\alpha}$ gets quotiented down to a point $[O_{\alpha}]\in\mathfrak{M}_{1}({\Lambda})$ with no stabilizer coming from this action.
@@
    If $y=0$ and $x\neq 0$, then $K$ acts freely and transitively on $O_{x}:=\{x\neq 0,y=0\}\subset X$. This orbit $O_{x}$ gets quotiented down to a point $[O_{x}]\in\mathfrak{M}_{1}({\Lambda})$ with no stabilizer coming from this action. Similarly, if $x=0$ and $y\neq 0$, then $K$ acts freely and transitively on $O_{y}:=\{x=0,y\neq 0\}\subset X$ and $O_{y}$ gets quotiented down to a point $[O_{y}]\in\mathfrak{M}_{1}({\Lambda})$ with no stabilizer coming from this action.
@@
    The point $O_{0}:=\{(0,0)\}\in X$ is a fixed points of the action. It is thus represented as a point $[O_{0}]\in\mathfrak{M}_{1}({\Lambda})$ with stabilizer $K$ coming from this action.
As before, in all the three cases above, there is always the additional stabilizer $H$ in all points of $\mathfrak{M}_{1}({\Lambda})$ coming for $BH$. Therefore, the stabilizers in $\mathfrak{M}^{(1)}_{\Lambda}=[X/G]$ of the points $O_{\alpha}$, $\alpha\neq 0$, $O_{x}$ and $O_{y}$ are $H$ and the stabilizer of $O_{0}$ is $G$. Now, the Hopf link admits the following three exact Lagrangian fillings:
@@
    An (unobstructed) immersed Lagrangian union $L_{0}$ of two embedded Lagrangian disks, with a unique intersection point. Each embedded Lagrangian disk fills a max-tb unknotted component of the Hopf link $\Lambda$.
@@
    The Lagrangian filling $L$ is encoded by the (reduced) substack
    $\left(\bigcup\nolimits_{\alpha\neq 0,-1}[O_{\alpha}]\right)\cup[O_{x}]\subset\mathfrak{M}_{1}({\Lambda})$
    where each point in this substack is a choice of $\operatorname{GL}_{1}$-local system in $L$. Similarly, the Lagrangian filling $L^{\prime}$ is encoded by the (reduced) substack
    $\left(\bigcup\nolimits_{\alpha\neq 0,-1}[O_{\alpha}]\right)\cup[O_{y}]\subset\mathfrak{M}_{1}({\Lambda}),$
    where again each point in this substack is a choice of $\operatorname{GL}_{1}$-local system in $L^{\prime}$. Note that for a choice of commutative ring $k$, both these substacks are isomorphic to $k^{*}/k^{*}$, which are indeed $\operatorname{GL}_{1}(k)$-local systems on a cylinder.
@@
    If one considers local systems on $L_{0}$, then $L_{0}$ is represented by the point $[O_{0}]\in\mathfrak{M}_{1}({\Lambda})$. [^5] In the appropriate sense, one could consider more general sheaves in $L_{0}$ so that this filling would be represented by the entirety of $\mathfrak{M}_{1}({\Lambda})$. That would be the case if we considered constructible sheaves on one of the disks stratified by the intersection point with the other disk and its complement. The fact that such sheaves give the entire $\mathfrak{M}_{1}({\Lambda})$ can be justified using Section 3 below, cf. Theorem 3.11. It is relevant to note that $[O_{0}]$ cannot be separated from either $[O_{x}]$ or $[O_{y}]$. We interpret this non-separatedness as an algebraic incarnation of the fact that both fillings $L$ and $L^{\prime}$, endowed with local systems giving $[O_{x}]$ and $[O_{y}]$, can be obtained from $L_{0}$ by resolving the immersed double point via Polterovich surgery. Therefore, in this sense, $[O_{0}]$ should lie as close as possible to both $[O_{x}]$ and $[O_{y}]$ inside of the moduli $\mathfrak{M}_{1}({\Lambda})$. $\Box$

### v2

@@ 2.3.1. Multiple basepoints per component
For the case of $\mathfrak{t}$ with multiple basepoints in one component of $\Lambda$, we can reduce to the case of one basepoint per component as follows. Consider a subset $\mathfrak{t}^{\prime}\subset\mathfrak{t}\subset\Lambda$ that still has at least one point per component. The homotopy colimit diagram (2.2) can be broken down into the following diagram
$\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)$$\prod_{i=1}^{|\mathfrak{t}|-|\mathfrak{t}^{\prime}|}\operatorname{Perf}(k)\times\operatorname{Perf}(k)$$\operatorname{Perf}(k)$$\operatorname{Sh}_{\Lambda}^{c}(\mathbb{R}^{2})_{0}$$\operatorname{Sh}_{\Lambda,\mathfrak{t}^{\prime}}^{c}(\mathbb{R}^{2})_{0}$$\operatorname{Sh}_{\Lambda,\mathfrak{t}}^{c}(\mathbb{R}^{2})_{0}.$}
@@
$\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})$$\mathfrak{M}_{n}({\Lambda,\mathfrak{t}^{\prime}})$$\mathfrak{M}_{n}({\Lambda})$$\operatorname{BGL}_{n}$$\prod_{i=1}^{|\mathfrak{t}|-|\mathfrak{t}^{\prime}|}\operatorname{BGL}_{n}\times\operatorname{BGL}_{n}$$\prod_{i=1}^{|\mathfrak{t}|}\operatorname{BGL}_{n}.$}
@@ 2.4. Microlocal rank 1 case of the basepoint action
Let us focus on $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\subset\mathfrak{M}({\Lambda,\mathfrak{t}})$, the component of $\mathfrak{M}({\Lambda,\mathfrak{t}})$ consisting of sheaves singularly supported in $(\Lambda,\mathfrak{t})$ with microlocal rank $1$. This is the component that has been most studied in the literature and it is of particular interest due to its connection to cluster algebras, see e.g. [CG22, CW24, CZ22, CGG+24, CGG+21, CGG+25]. The component $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ parametrizes sheaves with singular support on $\Lambda$ such that their microstalks at the basepoints of $\mathfrak{t}$ are 1-dimensional $k$-modules, all commonly identified. Similarly, we denote by $\mathfrak{M}_{1}({\Lambda})$ the component of $\mathfrak{M}({\Lambda})$ consisting of microlocal rank $1$ sheaves singularly supported in $\Lambda$.
@@
$\mathfrak{M}(\operatorname{Perf}(k))\quad\mbox{and}\quad\displaystyle\mathfrak{M}\left(\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)\right).$
@@
[…]
@@ 2.5. A few examples
Here are descriptions of the derived stacks $\mathfrak{M}_{1}({\Lambda})$ and $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ and the basepoint action in some cases of interest. In all the following examples, $\Lambda$ is a Legendrian (1)(-1)-closure of a positive braid, and by a direct generalization of [CNS19, Prop. 6.1], for any pseudo-perfect object in the category of sheaves valued in a simplicial commutative ring, the cotangent complexes of $\mathfrak{M}_{1}({\Lambda})$ and $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ are concentrated in non-negative degrees. Hence by [TV08, Theorem 2.2.2.6, Corollary 2.2.4.6] and [PRI12, Proposition 1.38] the derived stacks $\mathfrak{M}_{1}({\Lambda})$ and $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ are isomorphic to (underived) Artin stacks which parametrize microlocal rank 1 sheaves.
@@ 2.5.1. Legendrian knots
Let $\Lambda\subset(\mathbb{R}^{3},\xi_{\text{st}})\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{\text{st}})$ be a Legendrian knot with $\mathfrak{t}=\{t\}$ consisting of a single basepoint. Then $\Delta:H\longrightarrow G$ is the identity morphism and the homotopy pullback in Equation (2.6) is the base change along the identity morphism. Therefore $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\cong\mathfrak{M}_{1}({\Lambda})$ are isomorphic. Two remarks:
@@
    Let $\Lambda\subset(\mathbb{R}^{3},\xi_{\text{st}})$ be a Legendrian knots of the form $\Lambda=\Lambda_{\beta}$, $\beta$ a positive braid and $\Lambda_{\beta}$ its (1)(-1)-closure, see [CN22, Section 2.2]. Then the derived stacks $\mathfrak{M}_{1}({\Lambda_{\beta}})$ are isomorphic to quotient stacks $X_{\beta}/\operatorname{GL}_{1}]$ where $X_{\beta}$ are smooth affine varieties and $\operatorname{GL}_{1}$ acts trivially on $X_{\beta}$, cf. [CW24, Section 4.1]. In short, in this case we have
@@
    where $X_{\beta}$ is a smooth affine scheme. As stated in (ii) below, these $X_{\beta}$ are closely related to the braid varieties $X(\beta)$ studied in [CGG+24, CGG+21, CGG+25]. Also, equations expressing $X_{\beta}$ as a complete intersection can be readily read from a braid word for $\beta$, cf. [CGG+25, Section 3] (or [CN22, Section 5] from the perspective of the Legendrian contact dg algebra).
@@
    For a Legendrian knot with basepoints $\mathfrak{t}=\{t_{1},\ldots,t_{n}\}$, the action of $G/H$ on $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ is free and $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\cong\mathfrak{M}_{1}({\Lambda})\times\operatorname{GL}_{1}^{n-1}$, as per Equation (2.5). Let $\Lambda\subset(\mathbb{R}^{3},\xi_{\text{st}})$ be a Legendrian knot of the form $\Lambda=\Lambda_{\beta\delta(\beta)}$ , with one basepoint per strand of $\beta$. By [CW24, Section 4.1] or a direct generalization of [CL23, Proposition 5.8], we have
@@
    where $X(\beta)$ is the braid variety associated to $\beta$, see e.g. [CGG+24, Section 2] or [CGG+25, Section 3]. Specifically, per Equation (2.5), we have a regular isomorphism $X(\beta)\cong X_{\beta\delta(\beta)}\times\operatorname{GL}_{1}^{n-1}$, where $\delta(\beta)$ is a braid lift of the Demazure product of $\beta$ and $n$ is the number of strands.
@@ 2.5.2. Hopf link, Part I: $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})$ and $\mathfrak{M}_{1}({\Lambda})$
Let $\Lambda=\Lambda_{1}\cup\Lambda_{2}\subset(\mathbb{R}^{3},\xi_{\text{st}})\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{\text{st}})$ be the max-tb Hopf link, with both components $\Lambda_{1},\Lambda_{2}$ max-tb unknots. Consider the set of basepoints $\mathfrak{t}=\{t_{1},t_{2}\}$, with $t_{i}\in\Lambda_{i}$, one per component. The homotopy pullback in Equation (2.6) and the basepoint action are explicitly described as follows, cf. [CW24, Section 4.1]. Consider the polynomial ring $\mathbb{Z}[x,y]$ and its localization $\mathbb{Z}[x,y]_{(1+xy)}$ at the principal ideal $(1+xy)$. Define the affine scheme $X:=\operatorname{Spec}\mathbb{Z}[x,y]_{(1+xy)}$. One can show that the basepoint action from Lemma 2.3 can be described as follows: consider the algebraic action of $G:=\operatorname{GL}_{1}\times\operatorname{GL}_{1}$ on $X$ defined by
@@
Here we interpret the scheme $X$ as a 0-geometric Artin 0-stack, cf. [TV08, Prop. 2.1.2.1]. Note also that under $\Delta:H\longrightarrow G$ the diagonal morphism, the diagonal subgroup $\Delta(H)=\{(t_{1},t_{2}):t_{1}=t_{2}\}\subset G$ acts trivially on $X$. Then the moduli stacks of sheaves appearing in Equation (2.6) are
@@
The two input morphisms $\mathfrak{M}_{1}(m^{\ell}_{\Lambda,\mathfrak{t}}):[X/G]\longrightarrow BG$ and $B\Delta:BH\longrightarrow BG$ in the homotopy pullback Equation (2.6) are given by the canonical map $X/G]\longrightarrow BG$, induced by $X\longrightarrow\{\mbox{pt}\}$, and the diagonal map $B\Delta:\operatorname{BGL}_{1}\longrightarrow\operatorname{BGL}_{1}\times\operatorname{BGL}_{1}$. The induced morphism $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\longrightarrow BH$ is also given by the canonical map $X/H]\longrightarrow BH$. The last induced $\mathfrak{M}_{1}({\Lambda,\mathfrak{t}})\longrightarrow\mathfrak{M}_{1}({\Lambda})$ in Equation (2.6) is an algebraic quotient by $G/H$. This latter quotient can be realized by the action of $K:=\operatorname{GL}_{1}$ on $[X/H]$ induced by the action
@@
where we have taken the slice $\tau=t_{1}t_{2}^{-1}$ to represent the quotient $K\cong G/H$. In summary, with the notation and actions above, the homotopy pullback Equation (2.6) reads
[…]
where both vertical maps are the canonical projections induced by $X\longrightarrow\{\mbox{pt}\}$ and both horizontal maps can be modeled after $(G/H)$-principal bundles. In this context, the $(G/H)$-principal bundle $X/H]\longrightarrow[X/G]$ is classified by the composition $X/G]\longrightarrow BG\longrightarrow B(G/H)$.
@@ 2.5.3. Hopf link, Part II: Lagrangian fillings
The quotient stack $\mathfrak{M}_{1}({\Lambda})\cong[X/G]$ in Example 2.5.2 has a natural interpretation in terms of Lagrangian fillings of the Hopf link, as follows. There are three types of orbits in $X=\operatorname{Spec}\mathbb{Z}[x,y]_{(1+xy)}$ for the action Equation (2.7):
@@
    If $xy=\alpha$, $\alpha\in k$ and $\alpha\neq 0,-1$, then $K$ acts freely and transitively on $O_{\alpha}:=\{xy=\alpha\}\subset X$. Therefore, the orbit $O_{\alpha}$ gets quotiented down to a point $O_{\alpha}]\in\mathfrak{M}_{1}({\Lambda})$ with no stabilizer coming from this action.
@@
    If $y=0$ and $x\neq 0$, then $K$ acts freely and transitively on $O_{x}:=\{x\neq 0,y=0\}\subset X$. This orbit $O_{x}$ gets quotiented down to a point $O_{x}]\in\mathfrak{M}_{1}({\Lambda})$ with no stabilizer coming from this action. Similarly, if $x=0$ and $y\neq 0$, then $K$ acts freely and transitively on $O_{y}:=\{x=0,y\neq 0\}\subset X$ and $O_{y}$ gets quotiented down to a point $O_{y}]\in\mathfrak{M}_{1}({\Lambda})$ with no stabilizer coming from this action.
@@
    The point $O_{0}:=\{(0,0)\}\in X$ is a fixed points of the action. It is thus represented as a point $O_{0}]\in\mathfrak{M}_{1}({\Lambda})$ with stabilizer $K$ coming from this action.
As before, in all the three cases above, there is always the additional stabilizer $H$ in all points of $\mathfrak{M}_{1}({\Lambda})$ coming for $BH$. Therefore, the stabilizers in $\mathfrak{M}_{1}({\Lambda})=[X/G]$ of the points $O_{\alpha}$, $\alpha\neq 0$, $O_{x}$ and $O_{y}$ are $H$ and the stabilizer of $O_{0}$ is $G$. Now, the Hopf link admits the following three exact Lagrangian fillings:
@@
    An (unobstructed) immersed Lagrangian union $L_{0}$ of two embedded Lagrangian disks, with a unique transverse intersection point. Each embedded Lagrangian disk fills a max-tb unknotted component of the Hopf link $\Lambda$.
@@
    The Lagrangian filling $L$ is encoded by the (reduced) substack (⋃_α≠0,-1 [O_α])∪[O_x]⊂M_1(Λ) where each point in this substack is a choice of $\operatorname{GL}_{1}$-local system in $L$. Similarly, the Lagrangian filling $L^{\prime}$ is encoded by the (reduced) substack (⋃_α≠0,-1 [O_α])∪[O_y]⊂M_1(Λ), where again each point in this substack is a choice of $\operatorname{GL}_{1}$-local system in $L^{\prime}$. Note that for a choice of commutative ring $k$, both these substacks are isomorphic to $k^{*}/k^{*}$, which are indeed $\operatorname{GL}_{1}(k)$-local systems on a cylinder; see e.g. [TZ18, Section 2].
@@
    If one considers local systems on $L_{0}$, then $L_{0}$ is represented by the point $O_{0}]\in\mathfrak{M}_{1}({\Lambda})$. [^6] In the appropriate sense, one could consider more general sheaves in $L_{0}$ so that this filling would be represented by the entirety of $\mathfrak{M}_{1}({\Lambda})$. That would be the case if we considered constructible sheaves on one of the disks stratified by the intersection point with the other disk and its complement. The fact that such sheaves give the entire $\mathfrak{M}_{1}({\Lambda})$ can be justified using Section 3 below, cf. Theorem 3.13. It is relevant to note that $O_{0}]$ cannot be separated from either $O_{x}]$ or $O_{y}]$. We interpret this non-separatedness as an algebraic incarnation of the fact that both fillings $L$ and $L^{\prime}$, endowed with local systems giving $O_{x}]$ and $O_{y}]$, can be obtained from $L_{0}$ by resolving the immersed double point via Polterovich surgery. Therefore, in this sense, $O_{0}]$ should lie as close as possible to both $O_{x}]$ and $O_{y}]$ inside of the moduli $\mathfrak{M}_{1}({\Lambda})$. $\Box$